[10] 7. Use the Annihilator method to find a particular solution of the equation y – 4y + 4y = 2e24
Answer #1

Question 14 = f(t) (say) yil_47+4y= 2e2x The auxiliary equation is, 47+4= 0 + (-2)=0 Y = 2,2 The complementary function is,